marys avrage in 3 tests is 55.in 4 tests is 65.after 2 other tests her average mark is 55.find the average in the 2 other tests.
Answer by lwsshak3(6491)

marys avrage in 3 tests is 55.in 4 tests is 65.after 2 other tests her average mark is 55.find the average in the 2 other tests.
**
let x=average in the 2 other tests.
test scores total=3*55+4*65+2x=425+2x
number of tests=3+4+2=9
test scores total/number of tests=55
(425+2x)/9=55
(425+2x)=9*55=495
2x=495-425=70
x=70/2=35